Chris has invested $2,300 in two small businesses. Last year, the greeting card business made a 14% profit on his investment in the business, while the bicycle repair business showed a 6% loss. If Chris' net profit from the two businesses was $102, her investment in the greeting card business was $___?
Answer by checkley77(12569)

.14X-.06(2300-X)=102
.14X-138+.06X=102
.2X=102+138
.2X=240
X=240/.2
X=1200 INVESTED @ 14%
2300-1200=1100 INVESTED @ 6%
PROOF
.14*1200-.06*1100=102
168-66=102
102=102